A cheetah is traveling across the African countryside, initially moving at a constant velocity of 8.0 m/s. It sees a gazelle at t = 0 seconds and begins to accelerate in a straight line. At t = 4.0 seconds it has a velocity of 23 m/s and then continues to move at this constant velocity. At t = 5.0 seconds, the cheetah is standing still, having knocked over the gazelle.(a) What is the average acceleration of the cheetah from t = 0 to t = 5.0 seconds?
1
Expert's answer
2010-11-14T17:51:06-0500
v = v0 + at = 8 + 4*a = 23
a = 3.75 m/s^2
From t=4s to 5s the acceleration is equal to zero, thus the average is:
int(a(t)dt) / T = 4*3.75 / 5 = 3 m/s2